### v1

@@ 2.1 Exact and stable recovery
    (11) $\gamma\big(\alpha s-\alpha\sqrt{s}+(1-\alpha)s^{p}$ $H_{s,p}\big)x_{\min}+(1-\gamma)\|\bm{x}\|_{2}^{2}\leq f_{\alpha,\gamma}(\bm{x})$
    $\leq\gamma(\sqrt{s}-\alpha)\|\bm{x}\|_{2}+(1-\gamma)\|\bm{x}\|_{2}^{2}.$
@@
    If $\bm{x}$ has sparsity $s$ and $\|\bm{x}\|_{2}=r$, then $f_{\alpha,\gamma}(\bm{x})>0$ and $f_{\alpha,\gamma}(\bm{x})$ attains its minimum when $s=1$.
Proof 2.8.
It follows from $\alpha,\gamma\in(0,1)$ that the statements (a) and (b) can be readily verified from Lemma 2.4. We now prove the statement (c). Since $\alpha\in(0,1)$ and $\gamma\in(0,1)$, it follows that $s-\alpha\sqrt{s}\geq 0$ and so $f_{\alpha,\gamma}(\bm{x})\geq\gamma(s-\alpha\sqrt{s})x_{\min}+(1-\gamma)\|\bm{x}\|_{2}^{2}>0$, which shows that $f_{\alpha,\gamma}(\bm{x})>0$ for all $\bm{x}\neq 0$. Since $\|\bm{x}\|_{2}=r$, one has $f_{\alpha,\gamma}(\bm{x})=\gamma\|\bm{x}\|_{1}-\gamma\alpha r+(1-\gamma)r^{2}=\gamma(\|\bm{x}\|_{1}-\alpha r)+(1-\gamma)r^{2}$ and so the minimum of $\|\bm{x}\|_{1}$ is achieved when the support size satisfies $|\mathrm{supp}(\bm{x})|=s=1$. In this case, the vector $\bm{x}$ has only one nonzero entry of magnitude $r$, so $\|\bm{x}\|_{1}=r$. Hence, $f_{\alpha,\gamma}(\bm{x})$ attains its minimum when $s=1$.
Building upon the RIP-based analysis framework in [yin2015minimization, candes2005error], which established sufficient conditions for the exact recovery of the $\ell_{1}-\ell_{2}$ minimization and BP problem, we extend such a framework to CDCEN under the distinct condition.
Theorem 2.9.
Let $\alpha\in(0,1)$, $\gamma\in(0,1)$, and $p\geq 0$. Suppose that all the assumptions of Lemma 2.4 hold with $C=(3s)^{p}$. Let $\bm{\bar{x}}\in\mathbb{R}^{n}\setminus\{\bm{0}\}$ be an $s$-sparse signal, i.e., $\|\bm{\bar{x}}\|_{0}=s$ and $\|\bm{\bar{x}}\|_{2}\leq M(M>0)$, such that
$a(s,\gamma,\alpha,p)=\left(\frac{\gamma(\alpha\sqrt{3s}-\alpha+(1-\alpha)(3s)^{p}H_{3s,p}/\sqrt{3s})}{\gamma\sqrt{s}+\gamma\alpha+2(1-\gamma)M}\right)^{2}>1.$
If the sensing matrix $\bm{A}$ satisfies the RIP condition $\delta_{3s}+a(s,\gamma,\alpha,p)\,\delta_{4s}<a(s,\gamma,\alpha,p)-1$ and $\bm{A}\bm{\bar{x}}=\bm{b}$, then $\bm{\bar{x}}$ is a unique solution to the optimization problem (1).
Proof 2.10.
Let $\bm{x}$ be any feasible solution of (1) satisfying the linear constraint $\bm{Ax}=\bm{b}$ and has a smaller or equal function value, i.e., $\gamma(\|\bm{x}\|_{1}-\alpha\|\bm{x}\|_{2})+(1-\gamma)\|\bm{x}\|_{2}^{2}\leq\gamma(\|\bar{\bm{x}}\|_{1}-\alpha\|\bar{\bm{x}}\|_{2})+(1-\gamma)\|\bar{\bm{x}}\|_{2}^{2}$. Decomposing $\bm{x}$ as $\bm{x}=\bar{\bm{x}}+\bm{h}$ with $\bm{h}\in\ker(\bm{A})$, we show that $\bm{h}=\bm{0}$. To this end, let $\Lambda=\operatorname{supp}(\bar{\bm{x}})$ denote the support set of $\bar{\bm{x}}$, and decompose $\bm{h}$ into $\bm{h}=\bm{h}_{\Lambda}+\bm{h}_{\Lambda^{c}}$, where $\Lambda^{c}$ denotes the complement of $\Lambda$. Then
(12) $\gamma(\|\bm{x}\|_{1}-\alpha\|\bm{x}\|_{2})+(1-\gamma)\|\bm{x}\|_{2}^{2}\leq\gamma(\|\bar{\bm{x}}\|_{1}-\alpha\|\bar{\bm{x}}\|_{2})+(1-\gamma)\|\bar{\bm{x}}\|_{2}^{2}.$
@@
(13) $\|\bar{\bm{x}}+\bm{h}_{\Lambda}+\bm{h}_{\Lambda^{c}}\|_{1}=\|\bar{\bm{x}}+\bm{h}_{\Lambda}\|_{1}+\|\bm{h}_{\Lambda^{c}}\|_{1}\geq\|\bar{\bm{x}}\|_{1}-\|\bm{h}_{\Lambda}\|_{1}+\|\bm{h}_{\Lambda^{c}}\|_{1}.$
@@
(14) $\alpha\|\bar{\bm{x}}+\bm{h}_{\Lambda}+\bm{h}_{\Lambda^{c}}\|_{2}\leq\alpha(\|\bar{\bm{x}}\|_{2}+\|\bm{h}_{\Lambda}\|_{2}+\|\bm{h}_{\Lambda^{c}}\|_{2}).$
@@
(15) $(1-\gamma)\|\bar{\bm{x}}+\bm{h}_{\Lambda}+\bm{h}_{\Lambda^{c}}\|_{2}^{2}=(1-\gamma)(\|\bar{\bm{x}}+\bm{h}_{\Lambda}\|_{2}^{2}+\|\bm{h}_{\Lambda^{c}}\|_{2}^{2}).$
@@
(16) $\gamma(\|\bar{\bm{x}}+\bm{h}_{\Lambda}+\bm{h}_{\Lambda^{c}}\|_{1}-\alpha\|\bar{\bm{x}}+\bm{h}_{\Lambda}+\bm{h}_{\Lambda^{c}}\|_{2})+(1-\gamma)\|\bar{\bm{x}}+\bm{h}_{\Lambda}+\bm{h}_{\Lambda^{c}}\|_{2}^{2}$
$\geq\gamma\left[\|\bar{\bm{x}}\|_{1}-\|\bm{h}_{\Lambda}\|_{1}+\|\bm{h}_{\Lambda^{c}}\|_{1}-\alpha(\|\bar{\bm{x}}\|_{2}+\|\bm{h}_{\Lambda}\|_{2}+\|\bm{h}_{\Lambda^{c}}\|_{2})\right]$
$\quad\quad\quad\quad\quad\quad\quad\quad\;\;+(1-\gamma)(\|\bar{\bm{x}}+\bm{h}_{\Lambda}\|_{2}^{2}+\|\bm{h}_{\Lambda^{c}}\|_{2}^{2}).$
@@
(17) $\gamma\bigl[-\|\bm{h}_{\Lambda}\|_{1}$ $+\|\bm{h}_{\Lambda^{c}}\|_{1}-\alpha(\|\bm{h}_{\Lambda}\|_{2}+\|\bm{h}_{\Lambda^{c}}\|_{2})\bigr]+(1-\gamma)(\|\bar{\bm{x}}+\bm{h}_{\Lambda}\|_{2}^{2}+\|\bm{h}_{\Lambda^{c}}\|_{2}^{2})$
@@
$\quad\quad\quad\quad\quad\quad+(1-\gamma)\bigl(\|\bar{\bm{x}}\|_{2}^{2}+\|\bm{h}_{\Lambda^{c}}\|_{2}^{2}-2\|\bar{\bm{x}}\|_{2}\|\bm{h}_{\Lambda^{c}}\|_{2}\bigr).$
@@
(18) $\|\bm{h}_{\Lambda^{c}}\|_{1}-\alpha\|\bm{h}_{\Lambda^{c}}\|_{2}+\frac{1-\gamma}{\gamma}(\|\bm{h}_{\Lambda^{c}}\|_{2}^{2}+\|\bm{h}_{\Lambda}\|_{2}^{2}-2\|\bar{\bm{x}}\|_{2}\|\bm{h}_{\Lambda}\|_{2})\leq\|\bm{h}_{\Lambda}\|_{1}+\alpha\|\bm{h}_{\Lambda}\|_{2}.$
We sort the components of $\Lambda^{c}$ in descending order by magnitude and divide them into disjoint subsets $\Lambda_{i}^{c}$ with cardinality $3s$ (the last one possibly smaller), i.e., $\Lambda^{c}=\bigcup_{i=1}^{\ell}\Lambda_{i}^{c}$ and $\Lambda_{i}^{c}\cap\Lambda_{j}^{c}=\emptyset$ for $i\neq j$. Let $\Lambda_{0}^{c}=\Lambda\cup\Lambda_{1}^{c}$. Using the RIP property of $\bm{A}$ and $\bm{A}\bm{h}=0$, we have $0=\|\bm{Ah}\|_{2}\geq\left\|\bm{A}_{\Lambda_{0}^{c}}\bm{h}_{\Lambda_{0}^{c}}\right\|_{2}-\left\|\sum_{i=2}^{\ell}\bm{A}_{\Lambda_{i}^{c}}\bm{h}_{\Lambda_{i}^{c}}\right\|_{2}\geq\left\|\bm{A}_{\Lambda_{0}^{c}}\bm{h}_{\Lambda_{0}^{c}}\right\|_{2}-\sum_{i=2}^{\ell}\left\|\bm{A}_{\Lambda_{i}^{c}}\bm{h}_{\Lambda_{i}^{c}}\right\|_{2}$. Set $A\coloneqq\sqrt{1-\delta_{4s}}$ and $B\coloneqq\sqrt{1+\delta_{3s}}$. It follows from the RIP condition that $0=\|\bm{Ah}\|_{2}\geq A\|\bm{h}_{\Lambda_{0}^{c}}\|_{2}-B\sum_{i=2}^{\ell}\|\bm{h}_{\Lambda_{i}^{c}}\|_{2}$. By the statement (b) in Lemma 2.7, for any $i\geq 2$ and $t\in\Lambda_{i}^{c}$, one has
@@
$\leq\min_{r\in\Lambda_{i-1}^{c}}|h_{r}|+\frac{(1-\gamma)\|\bm{h}_{\Lambda_{i-1}^{c}}\|_{2}^{2}}{\gamma(3\alpha s-\alpha\sqrt{3s}+(1-\alpha)3^{p}s^{p}H_{3s,p})}$
@@
(19) $\sum_{i=2}^{\ell}\|\bm{h}_{\Lambda_{i}^{c}}\|_{2}$ $\leq\sqrt{3s}\sum_{i=1}^{\ell-1}\frac{\gamma\left(\|\bm{h}_{\Lambda_{i}^{c}}\|_{1}-\alpha\|\bm{h}_{\Lambda_{i}^{c}}\|_{2}\right)+(1-\gamma)\|\bm{h}_{\Lambda_{i}^{c}}\|_{2}^{2}}{\gamma\left(3\alpha s-\alpha\sqrt{3s}+(1-\alpha)\,3^{p}s^{p}H_{3s,p}\right)}$
@@
$\leq\sqrt{3s}\cdot\frac{\gamma\sum_{i=1}^{\ell}\|\bm{h}_{\Lambda_{i}^{c}}\|_{1}-\gamma\alpha\sum_{i=1}^{\ell}\|\bm{h}_{\Lambda_{i}^{c}}\|_{2}+(1-\gamma)\sum_{i=1}^{\ell}\|\bm{h}_{\Lambda_{i}^{c}}\|_{2}^{2}}{\gamma\left(3\alpha s-\alpha\sqrt{3s}+(1-\alpha)\,3^{p}s^{p}H_{3s,p}\right)}.$
@@
By $\sum_{i=1}^{\ell}\|\bm{h}_{\Lambda_{i}^{c}}\|_{1}=\|\bm{h}_{\Lambda^{c}}\|_{1}$, $(1-\gamma)\sum_{i=1}^{\ell}\|\bm{h}_{\Lambda_{i}^{c}}\|_{2}^{2}=(1-\gamma)\|\bm{h}_{\Lambda^{c}}\|_{2}^{2}$, and the inequality $\gamma\alpha\sum_{i=1}^{\ell}\|\bm{h}_{\Lambda_{i}^{c}}\|_{2}\geq\gamma\alpha\sqrt{\sum_{i=1}^{\ell}\|\bm{h}_{\Lambda_{i}^{c}}\|_{2}^{2}}=\gamma\alpha\|\bm{h}_{\Lambda^{c}}\|_{2}$, and substitute (18) into (19) to obtain
$\sum_{i=2}^{\ell}\|\bm{h}_{\Lambda_{i}^{c}}\|_{2}$ $\overset{a}{\leq}\sqrt{3s}\cdot\frac{\left(\gamma\sqrt{s}+\gamma\alpha\right)\|\bm{h}_{\Lambda}\|_{2}+(\gamma-1)\left(\|\bm{h}_{\Lambda}\|_{2}^{2}-2\|\bm{h}_{\Lambda}\|_{2}\|\bm{\bar{x}}\|_{2}\right)}{\gamma\left(3\alpha s-\alpha\sqrt{3s}+(1-\alpha)3^{p}s^{p}H_{3s,p}\right)}$
@@
$\overset{b}{\leq}\frac{\left(\gamma\sqrt{s}+\gamma\alpha+2(1-\gamma)M\right)\|\bm{h}_{\Lambda}\|_{2}}{\gamma\left(\alpha\sqrt{3s}-\alpha+(1-\alpha)3^{p}s^{p}H_{3s,p}/\sqrt{3s}\right)},$
@@
where inequality $\overset{a}{\leq}$ holds due to condition $\gamma\|\bm{h}_{\Lambda}\|_{1}\leq\gamma\sqrt{s}\|\bm{h}_{\Lambda}\|_{2}$ and inequality $\overset{b}{\leq}$ follows from the boundedness of the signal $\bm{\bar{x}}$, i.e., $\|\bm{\bar{x}}\|_{2}\leq M$. Let
$\sqrt{a(s,\gamma,\alpha,p)}=\frac{\gamma(\alpha\sqrt{3s}-\alpha+(1-\alpha)(3s)^{p}H_{3s,p}/\sqrt{3s})}{\gamma\sqrt{s}+\gamma\alpha+2(1-\gamma)M}.$
@@
Then $\sum_{i=2}^{\ell}\|\bm{h}_{\Lambda_{i}^{c}}\|_{2}\leq\frac{\|\bm{h}_{\Lambda}\|_{2}}{\sqrt{a(s,\gamma,\alpha,p)}}$. It follows that $0\geq A\|\bm{h}_{\Lambda_{0}^{c}}\|_{2}-B\sum_{i=2}^{\ell}\|\bm{h}_{\Lambda_{i}^{c}}\|_{2}\geq A\|\bm{h}_{\Lambda_{0}^{c}}\|_{2}-\frac{B}{\sqrt{a(s,\lambda,\alpha,p)}}\|\bm{h}_{\Lambda}\|_{2}\overset{a}{\geq}\left(A-\frac{B}{\sqrt{a(s,\lambda,\alpha,p)}}\right)\|\bm{h}_{\Lambda_{0}^{c}}\|_{2}$, where the inequality $\overset{a}{\geq}$ holds because $\|\bm{h}_{\Lambda_{0}^{c}}\|_{2}^{2}=\|\bm{h}_{\Lambda}+\bm{h}_{\Lambda_{1}^{c}}\|_{2}^{2}\geq\|\bm{h}_{\Lambda}\|_{2}^{2}$. Since $a(s,\gamma,\alpha,p)>1$ and $\delta_{3s}+a(s,\gamma,\alpha,p)\delta_{4s}<a(s,\gamma,\alpha,p)-1$, it follows that $\sqrt{1-\delta_{4s}}-\frac{\sqrt{1+\delta_{3s}}}{\sqrt{a(s,\lambda,\alpha,p)}}>0$, which implies $\bm{h}_{\Lambda_{0}^{c}}=\bm{0}$ and so $\bm{h}=\bm{0}$. This proves the uniqueness of $\bar{\bm{x}}$.
@@
Theorem 2.9 establishes a RIP–based sufficient condition ensuring the uniqueness of the solution recovered by the measure $\gamma\left(\|\bm{x}\|_{1}-\alpha\|\bm{x}\|_{2}\right)+\left(1-\gamma\right)\|\bm{x}\|_{2}^{2}$. This result parallels the classical RIP condition for exact recovery in BP [candes2005error], yet it is specifically tailored to the composite nonconvex objective above. Unlike the $\ell_{1}$ formulation in BP, this condition explicitly characterizes the interplay between the sparsity-inducing term $\|\bm{x}\|_{1}-\alpha\|\bm{x}\|_{2}$ and the quadratic regularization $\|\bm{x}\|_{2}^{2}$.
@@
At the end of this subsection, we extend the stable recovery property of [yin2015minimization] to demonstrate the stable recovery property of CDCEN in the presence of measurement noise.
Theorem 2.11.
@@
Let $\alpha\in(0,1)$ and $\gamma\in(0,1)$, and suppose that all the conditions in Theorem 2.9 are satisfied with $\bm{A}\bm{\bar{x}}+\bm{e}=\bm{b}$, where $\bm{e}\in\mathbb{R}^{m}$ denotes an arbitrary noise with $\|\bm{e}\|_{2}\leq\tau$. Let $\bm{x}^{\mathrm{opt}}$ be an optimal solution of the noise-aware problem
$\min_{\bm{x}\in\mathbb{R}^{n}}\gamma\big(\|\bm{x}\|_{1}-\alpha\|\bm{x}\|_{2}\big)+(1-\gamma)\|\bm{x}\|_{2}^{2}\quad\text{subject to}\quad\|\bm{A}\bm{x}-\bm{b}\|_{2}\leq\tau.$

### v2

@@ II-A Exact and stable recovery
    $\gamma\big(\alpha s-\alpha\sqrt{s}+(1-\alpha)s^{p}H_{s,p}\big)x_{\min}+(1-\gamma)\|\bm{x}\|_{2}^{2}$ (11)
    $\leq r(\bm{x})\leq\gamma(\sqrt{s}-\alpha)\|\bm{x}\|_{2}+(1-\gamma)\|\bm{x}\|_{2}^{2}.$
@@
    If $\bm{x}$ has sparsity $s$ and $\|\bm{x}\|_{2}=r$, then $r(\bm{x})>0$ and $r(\bm{x})$ attains its minimum when $s=1$.
Proof.
It follows from $\alpha,\gamma\in(0,1)$ that the statements (a) and (b) can be readily verified from Lemma 3. We now prove the statement (c). Since $\alpha\in(0,1)$ and $\gamma\in(0,1)$, it follows that $s-\alpha\sqrt{s}\geq 0$ and so $r(\bm{x})\geq\gamma(s-\alpha\sqrt{s})x_{\min}+(1-\gamma)\|\bm{x}\|_{2}^{2}>0$, which shows that $r(\bm{x})>0$ for all $\bm{x}\neq 0$. Since $\|\bm{x}\|_{2}=r$, one has $r(\bm{x})=\gamma\|\bm{x}\|_{1}-\gamma\alpha r+(1-\gamma)r^{2}=\gamma(\|\bm{x}\|_{1}-\alpha r)+(1-\gamma)r^{2}$ and so the minimum of $\|\bm{x}\|_{1}$ is achieved when the support size satisfies $|\mathrm{supp}(\bm{x})|=s=1$. In this case, the vector $\bm{x}$ has only one nonzero entry of magnitude $r$, so $\|\bm{x}\|_{1}=r$. Hence, $r(\bm{x})$ attains its minimum when $s=1$. ∎
Building upon the RIP-based analysis framework in [41, 6], which established sufficient conditions for the exact recovery of the $\ell_{1}-\ell_{2}$ minimization and BP problem, we extend such a framework to CDCEN under the distinct condition.
Theorem 6.
Let $\alpha\in(0,1)$, $\gamma\in(0,1)$, and $p\geq 0$. Suppose that all the assumptions of Lemma 3 hold with $C=(3s)^{p}$. Let $\bm{\bar{x}}\in\mathbb{R}^{n}\setminus\{\bm{0}\}$ be an $s$-sparse signal, i.e., $\|\bm{\bar{x}}\|_{0}=s$ and $\|\bm{\bar{x}}\|_{2}\leq R(R>0)$, such that
$a(s,\gamma,\alpha,p)=\left(\frac{\gamma(\alpha\sqrt{3s}-\alpha+p(\alpha,s))}{\gamma\sqrt{s}+\gamma\alpha+2(1-\gamma)R}\right)^{2}>1,$
where $p(\alpha,s)=(1-\alpha)(3s)^{p}H_{3s,p}/\sqrt{3s}$. If the sensing matrix $\bm{A}$ satisfies the RIP condition $\delta_{3s}+a(s,\gamma,\alpha,p)\,\delta_{4s}<a(s,\gamma,\alpha,p)-1$ and $\bm{A}\bm{\bar{x}}=\bm{b}$, then $\bm{\bar{x}}$ is a unique solution to the optimization problem (1).
Proof.
Let $\bm{x}$ be any feasible solution of (1) satisfying the linear constraint $\bm{Ax}=\bm{b}$ and has a smaller or equal function value, i.e., $r(\bm{x})\leq r(\bar{\bm{x}})$. Decomposing $\bm{x}$ as $\bm{x}=\bar{\bm{x}}+\bm{h}$ with $\bm{h}\in\ker(\bm{A})$, we show that $\bm{h}=\bm{0}$. To this end, let $\Lambda=\operatorname{supp}(\bar{\bm{x}})$ denote the support set of $\bar{\bm{x}}$, and decompose $\bm{h}$ into $\bm{h}=\bm{h}_{\Lambda}+\bm{h}_{\Lambda^{c}}$, where $\Lambda^{c}$ denotes the complement of $\Lambda$. Then
$\gamma(\|\bm{x}\|_{1}$ $-\alpha\|\bm{x}\|_{2})+(1-\gamma)\|\bm{x}\|_{2}^{2}$ (12)
$\leq\gamma(\|\bar{\bm{x}}\|_{1}-\alpha\|\bar{\bm{x}}\|_{2})+(1-\gamma)\|\bar{\bm{x}}\|_{2}^{2}.$
@@
$\|\bar{\bm{x}}+\bm{h}_{\Lambda}+\bm{h}_{\Lambda^{c}}\|_{1}$ $=\|\bar{\bm{x}}+\bm{h}_{\Lambda}\|_{1}+\|\bm{h}_{\Lambda^{c}}\|_{1}$ (13)
$\geq\|\bar{\bm{x}}\|_{1}-\|\bm{h}_{\Lambda}\|_{1}+\|\bm{h}_{\Lambda^{c}}\|_{1}.$
@@
$\alpha\|\bar{\bm{x}}+\bm{h}_{\Lambda}+\bm{h}_{\Lambda^{c}}\|_{2}\leq\alpha(\|\bar{\bm{x}}\|_{2}+\|\bm{h}_{\Lambda}\|_{2}+\|\bm{h}_{\Lambda^{c}}\|_{2}).$ (14)
@@
$(1-\gamma)\|\bar{\bm{x}}+\bm{h}_{\Lambda}+\bm{h}_{\Lambda^{c}}\|_{2}^{2}=(1-\gamma)(\|\bar{\bm{x}}+\bm{h}_{\Lambda}\|_{2}^{2}+\|\bm{h}_{\Lambda^{c}}\|_{2}^{2}).$ (15)
@@
$\gamma(\|\bar{\bm{x}}+\bm{h}_{\Lambda}+\bm{h}_{\Lambda^{c}}\|_{1}-\alpha\|\bar{\bm{x}}+\bm{h}_{\Lambda}+\bm{h}_{\Lambda^{c}}\|_{2})+$ (16)
$(1-\gamma)\|\bar{\bm{x}}+\bm{h}_{\Lambda}+\bm{h}_{\Lambda^{c}}\|_{2}^{2}\geq\gamma[\|\bar{\bm{x}}\|_{1}-\|\bm{h}_{\Lambda}\|_{1}$
$+\|\bm{h}_{\Lambda^{c}}\|_{1}-\alpha(\|\bar{\bm{x}}\|_{2}+\|\bm{h}_{\Lambda}\|_{2}+\|\bm{h}_{\Lambda^{c}}\|_{2})]$
$\quad\quad\quad\quad\;\;+(1-\gamma)(\|\bar{\bm{x}}+\bm{h}_{\Lambda}\|_{2}^{2}+\|\bm{h}_{\Lambda^{c}}\|_{2}^{2}).$
@@
$\gamma\bigl[$ $-\|\bm{h}_{\Lambda}\|_{1}+\|\bm{h}_{\Lambda^{c}}\|_{1}-\alpha(\|\bm{h}_{\Lambda}\|_{2}+\|\bm{h}_{\Lambda^{c}}\|_{2})\bigr]$ (17)
$+(1-\gamma)(\|\bar{\bm{x}}+\bm{h}_{\Lambda}\|_{2}^{2}+\|\bm{h}_{\Lambda^{c}}\|_{2}^{2})$
@@
$\quad+(1-\gamma)\bigl(\|\bar{\bm{x}}\|_{2}^{2}+\|\bm{h}_{\Lambda^{c}}\|_{2}^{2}-2\|\bar{\bm{x}}\|_{2}\|\bm{h}_{\Lambda^{c}}\|_{2}\bigr).$
@@
$\|\bm{h}_{\Lambda^{c}}\|_{1}$ $-\alpha\|\bm{h}_{\Lambda^{c}}\|_{2}+\frac{1-\gamma}{\gamma}(\|\bm{h}_{\Lambda^{c}}\|_{2}^{2}+\|\bm{h}_{\Lambda}\|_{2}^{2}$ (18)
$-2\|\bar{\bm{x}}\|_{2}\|\bm{h}_{\Lambda}\|_{2})\leq\|\bm{h}_{\Lambda}\|_{1}+\alpha\|\bm{h}_{\Lambda}\|_{2}.$
We sort the components of $\Lambda^{c}$ in descending order by magnitude and divide them into disjoint subsets $\Lambda_{i}^{c}$ with cardinality $3s$ (the last one possibly smaller), i.e., $\Lambda^{c}=\bigcup_{i=1}^{\ell}\Lambda_{i}^{c}$ and $\Lambda_{i}^{c}\cap\Lambda_{j}^{c}=\emptyset$ for $i\neq j$. Let $\Lambda_{0}^{c}=\Lambda\cup\Lambda_{1}^{c}$. Using the RIP property of $\bm{A}$ and $\bm{A}\bm{h}=0$, we have $0=\|\bm{Ah}\|_{2}\geq\left\|\bm{A}_{\Lambda_{0}^{c}}\bm{h}_{\Lambda_{0}^{c}}\right\|_{2}-\left\|\sum_{i=2}^{\ell}\bm{A}_{\Lambda_{i}^{c}}\bm{h}_{\Lambda_{i}^{c}}\right\|_{2}\geq\left\|\bm{A}_{\Lambda_{0}^{c}}\bm{h}_{\Lambda_{0}^{c}}\right\|_{2}-\sum_{i=2}^{\ell}\left\|\bm{A}_{\Lambda_{i}^{c}}\bm{h}_{\Lambda_{i}^{c}}\right\|_{2}$. Set $A\coloneqq\sqrt{1-\delta_{4s}}$ and $B\coloneqq\sqrt{1+\delta_{3s}}$. It follows from the RIP condition that $0=\|\bm{Ah}\|_{2}\geq A\|\bm{h}_{\Lambda_{0}^{c}}\|_{2}-B\sum_{i=2}^{\ell}\|\bm{h}_{\Lambda_{i}^{c}}\|_{2}$. By the statement (b) in Lemma 5, for any $i\geq 2$ and $t\in\Lambda_{i}^{c}$, one has
@@
$\sum_{i=2}^{\ell}\|\bm{h}_{\Lambda_{i}^{c}}\|_{2}$ (19)
@@
$\leq\sqrt{3s}\cdot\frac{\gamma\sum_{i=1}^{\ell}\|\bm{h}_{\Lambda_{i}^{c}}\|_{1}-\gamma\alpha\ell_{h^{c}}+(1-\gamma)\sum_{i=1}^{\ell}\|\bm{h}_{\Lambda_{i}^{c}}\|_{2}^{2}}{\gamma\left(3\alpha s-\alpha\sqrt{3s}+(1-\alpha)\,3^{p}s^{p}H_{3s,p}\right)},$
@@
where $\ell_{h^{c}}=\sum_{i=1}^{\ell}\|\bm{h}_{\Lambda_{i}^{c}}\|_{2}$. By $\sum_{i=1}^{\ell}\|\bm{h}_{\Lambda_{i}^{c}}\|_{1}=\|\bm{h}_{\Lambda^{c}}\|_{1}$, $(1-\gamma)\sum_{i=1}^{\ell}\|\bm{h}_{\Lambda_{i}^{c}}\|_{2}^{2}=(1-\gamma)\|\bm{h}_{\Lambda^{c}}\|_{2}^{2}$, and $\gamma\alpha\sum_{i=1}^{\ell}\|\bm{h}_{\Lambda_{i}^{c}}\|_{2}\geq\gamma\alpha\sqrt{\sum_{i=1}^{\ell}\|\bm{h}_{\Lambda_{i}^{c}}\|_{2}^{2}}=\gamma\alpha\|\bm{h}_{\Lambda^{c}}\|_{2}$, and substitute (18) into (19) to obtain
@@
$\sum_{i=2}^{\ell}\|\bm{h}_{\Lambda_{i}^{c}}\|_{2}\overset{b}{\leq}\frac{\left(\gamma\sqrt{s}+\gamma\alpha+2(1-\gamma)R\right)\|\bm{h}_{\Lambda}\|_{2}}{\gamma\left(\alpha\sqrt{3s}-\alpha+(1-\alpha)3^{p}s^{p}H_{3s,p}/\sqrt{3s}\right)},$
@@
where inequality $\overset{a}{\leq}$ holds due to condition $\gamma\|\bm{h}_{\Lambda}\|_{1}\leq\gamma\sqrt{s}\|\bm{h}_{\Lambda}\|_{2}$ and inequality $\overset{b}{\leq}$ follows from the boundedness of the signal $\bm{\bar{x}}$, i.e., $\|\bm{\bar{x}}\|_{2}\leq R$. Let
$\sqrt{a(s,\gamma,\alpha,p)}=\frac{\gamma(\alpha\sqrt{3s}-\alpha+(1-\alpha)(3s)^{p}H_{3s,p}/\sqrt{3s})}{\gamma\sqrt{s}+\gamma\alpha+2(1-\gamma)R}.$
@@
Then $\sum_{i=2}^{\ell}\|\bm{h}_{\Lambda_{i}^{c}}\|_{2}\leq\frac{\|\bm{h}_{\Lambda}\|_{2}}{\sqrt{a(s,\gamma,\alpha,p)}}$. It follows that $0\geq A\|\bm{h}_{\Lambda_{0}^{c}}\|_{2}-B\sum_{i=2}^{\ell}\|\bm{h}_{\Lambda_{i}^{c}}\|_{2}\geq A\|\bm{h}_{\Lambda_{0}^{c}}\|_{2}-\frac{B}{\sqrt{a(s,\gamma,\alpha,p)}}\|\bm{h}_{\Lambda}\|_{2}\overset{a}{\geq}\left(A-\frac{B}{\sqrt{a(s,\lambda,\alpha,p)}}\right)\|\bm{h}_{\Lambda_{0}^{c}}\|_{2}$, where the inequality $\overset{a}{\geq}$ holds because $\|\bm{h}_{\Lambda_{0}^{c}}\|_{2}^{2}=\|\bm{h}_{\Lambda}+\bm{h}_{\Lambda_{1}^{c}}\|_{2}^{2}\geq\|\bm{h}_{\Lambda}\|_{2}^{2}$. Since $a(s,\gamma,\alpha,p)>1$ and $\delta_{3s}+a(s,\gamma,\alpha,p)\delta_{4s}<a(s,\gamma,\alpha,p)-1$, it follows that $\sqrt{1-\delta_{4s}}-\frac{\sqrt{1+\delta_{3s}}}{\sqrt{a(s,\lambda,\alpha,p)}}>0$, which implies $\bm{h}_{\Lambda_{0}^{c}}=\bm{0}$ and so $\bm{h}=\bm{0}$. This proves the uniqueness of $\bar{\bm{x}}$. ∎
@@
Theorem 6 establishes a RIP–based sufficient condition ensuring the uniqueness of the solution recovered by the measure $\gamma\left(\|\bm{x}\|_{1}-\alpha\|\bm{x}\|_{2}\right)+\left(1-\gamma\right)\|\bm{x}\|_{2}^{2}$. This result parallels the classical RIP condition for exact recovery in BP [6], yet it is specifically tailored to the composite nonconvex objective above. Unlike the $\ell_{1}$ formulation in BP, this condition explicitly characterizes the interplay between the sparsity-inducing term $\|\bm{x}\|_{1}-\alpha\|\bm{x}\|_{2}$ and the quadratic regularization $\|\bm{x}\|_{2}^{2}$.
@@
At the end of this subsection, we extend the stable recovery property of [41] to demonstrate the stable recovery property of CDCEN in the presence of measurement noise.
Theorem 7.
@@
Let $\alpha\in(0,1)$ and $\gamma\in(0,1)$, and suppose that all the conditions in Theorem 6 are satisfied with $\bm{A}\bm{\bar{x}}+\bm{e}=\bm{b}$, where $\bm{e}\in\mathbb{R}^{m}$ denotes an arbitrary noise with $\|\bm{e}\|_{2}\leq\tau$. Let $\bm{x}^{\mathrm{opt}}$ be an optimal solution of the noise-aware problem
$\min_{\bm{x}\in\mathbb{R}^{n}}r(\bm{x})\quad\text{subject to}\quad\|\bm{A}\bm{x}-\bm{b}\|_{2}\leq\tau.$